OpenStudy (anonymous):

Evaluate the Integral from 1 to 2, (4+u^2)/u^3 du

OpenStudy (anonymous):

distribute the denominator between the two terms :) It should reduce to standard forms. Then integrate and apply the limits :)

OpenStudy (anonymous):

\[4 \int\limits_{2}^{1} \frac{ 1 }{ u } * \frac{ 1+u ^{2} }{ u ^{2}}\]

OpenStudy (anonymous):

Is that correct by any chance?

OpenStudy (anonymous):

Aha sorry I'm terrible at this integral stuff :(

OpenStudy (anonymous):

i meant to convert \[\int\limits_{1}^{2}\frac{ 4+u^{2} }{ u ^{3} } = \int\limits_{1}^{2} \frac{ 4 }{ u ^{3} } + \frac{ u ^{2} }{ u ^{3} }\]

OpenStudy (anonymous):

both those are standard integrals . you should be able to evaluate them easily :)

OpenStudy (anonymous):

Ahhhh okay that makes sense now...Thanks a bunch! :)

OpenStudy (anonymous):

you're welcome :)